# references.bib

### v1

@@ Appendix B Representational geometries across different deep neural network architectures
Note that the proof is a minor extension of Lemma S1.3 in [lippl_mathematical_2024]. We present it here in a self-contained manner. We consider a nonlinearity
@@
By prior work, [cho_kernel_2009, tsuchida_invariance_2018, tsuchida_richer_2019, han_fast_2022],
@@
B.2 Extended analysis of representational salience
To complement Section 4.2, we analyze the impact of different magnitudes for different components. In particular, we consider a disentangled representation whose first component has a salience between $s\in[0.1,0.5]$ and whose second component accordingly has a salience $1-s$ (Fig. 6a). As becomes deeper, the more salient component (in this case component 2) increasingly dominates the representation. While the salience of the full conjunction still eventually converges to one for most nonlinearities, it takes longer to do so for a less balanced input representation. Further, for a Gaussian nonlinearity, an imbalanced representation actually decreases the limit salience the representation appears to be converging to for the full conjunction.
In Fig. 6, we plot the different saliences for an input with four components. Now the salience of overlap 2 and 3 both first increase and then decrease and, just like for inputs with three components (Fig. 2), a Gaussian and rectified quadratic nonlinearity yields a particularly high salience for these intermediate conjunctions. Notably, they appear to be trading off the salience of these conjunctions differently: the rectified quadratic nonlinearity more strongly emphasizes overlaps of three whereas the Gaussian nonlinearity more strongly emphasizes overlaps of two.
@@ C.1 Models
We fit the kernel models by hand-specifying the kernel and fitting either a support vector regression or classification using scikit-learn [pedregosa_scikit-learn_2011].
@@
All networks were trained with Pytorch and Pytorch Lightning [paszke_pytorch_2019]. We consider ReLU networks with one hidden layer and $H=1000$ units. We initialize by $\sigma\sqrt{2/H}$, considering $\sigma\in[10^{-6},1]$. In particular, when reporting results on rich networks (without further specification), we assume $\sigma=10^{-6}$. When reporting results on lazy network, we assume $\sigma=1$.
@@
We considered networks with four convolutional layers (kernel size is five, two layers have 32 filters, two have 64 filters) and two densely connected layers (with 512 and 1024 units). Each layer is followed by a ReLU nonlinearity, and the convolutional stage is followed by a max pooling operation. All weights are initialized with He initialization [he_delving_2015].
@@
C.2 Reproducibility
Computational simulations.
We ran all computational simulations either on a CPU or on a single GPU. Each kernel experiment (total number: 53 experiments) ran within fifteen minutes, each rich network (total number: 450 experiments) and ConvNet experiment (total number: 600 experiments) took between one and two hours, and each ViT or ResNet experiment (total number: 140 experiments) took between five and ten hours. We used a computational cluster to run many of the experiments in parallel. As a rough estimate, the total experiments therefore required 900 hours of CPU computations and 2,600 hours of GPU computations.
Code repository.
The code required to reproduce all experiments can be found under https://github.com/sflippl/compositional-generalization. We additionally uploaded the repository including all simulated data to Zenodo (https://doi.org/10.5281/zenodo.11308156).
C.3 Additivity analysis
To analyze how well a conjunction-wise additive computation can describe network behavior, we considered as the set of possible features a concatenation of one-hot vectors coding for each possible conjunction. We then removed all features that are constant at zero on the training dataset and used linear regression to try and predict network behavior on both training and test set for all remaining features. The resulting $R^{2}$ defines the “additivity” of the network behavior (i.e. $R^{2}=1$ indicates full conjunction-wise additivity). Furthermore, we can use the inferred values assigned to these different conjunctions to compare kernel models, rich and lazy networks, and convolutional networks. Note that for the convolutional networks, we first average the model predictions across all different images instantiating a given compositional input.
Appendix D Compositional tasks
@@ D.1 Symbolic addition
Note that the sum over all dual coefficients is given by $2b+c$. Let $p:=|\mathcal{W}|$ and $q:=|\mathcal{V}|-|\mathcal{W}|$. Then, setting $\delta_{2}:=\kappa_{2}-\kappa_{0}$ and $\delta_{1}:=\kappa_{1}-\kappa_{0}$, the set of dual equations is given by
@@
D.1.2 Behavior of kernel models on other tasks
Figure 7: Kernel models on dispersed training sets. a, Four example training sets containing two random instances containing each item. b, Generalization loss as a function of the representation’s overlap salience. c,d, Inferred values for c) four example tasks and d) the average across fifty randomly drawn training sets.
To test whether the observed behavior was specific to the kinds of training sets investigated in Proposition 4.4 and the main text, we additionally considered dispersed, randomly generated training sets with two randomly drawn trials containing each item (Fig. 7a). We found that the generalized loss (averaged across all cases) increased roughly linearly with salience (Fig. 7b). For individual training sets, the inferred values were distorted in a complex, irregular manner (Fig. 7c). Further, because the training sets were no longer symmetric, the items took on different values when presented as the first or second component. On average, however, the inferred values were roughly proportional to their actual value and the factor of proportionality still decreased with salience (Fig. 7d).
D.1.3 Behavior of rich networks
Figure 8: Behavior of rich networks on symbolic addition. a, Additivity as a function of initialization scale in networks with one hidden layer. b, Similarity between distinct, identical, and overlapping trials, before and after training in the lazy or rich regime. c, Inferred values in the lazy and rich regime for networks with one hidden layer. d, Zoomed-in inferred values for interpolation part of the task involving interpolation and extrapolation. e, Additivity as a function of the number of hidden layers. f, Inferred values for networks with different numbers of layers.
We first trained rich networks with one hidden layer on the different variants of symbolic addition. We found that they generally were highly additive ($R^{2}>0.992$, Fig. 8a). Fig. 8b confirms that the ReLU networks indeed change their representation in the rich regime: at initialization, the similarity between different trials is approximately clustered by whether those trials are distinct, overlapping, or identical. After lazy training, this remains unchanged, whereas after rich training, the similarities look entirely different. Finally, the inferred values look similar on extrapolation, but do not exhibit a memorization leak on interpolation (Fig. 8c). On the task involving both interpolation and extrapolation, the networks suffer from a memorization leak on the extrapolaton region but not the interpolation region (Fig. 8d). This indicates that rich networks tend to perform better at interpolation, unlike lazy networks which are not sensitive to this difference.
D.1.4 Behavior of deep networks
Next, we trained networks with one to four hidden layers, using $\sigma=0.1$. We found that the networks were generally highly additive ($R^{2}>0.985$, Fig. 8e) and inferred highly similar values. In particular, deeper networks also suffered from a memorization leak on extrapolation but not interpolation.
D.1.5 Behavior of vision models trained on MNIST and CIFAR-10
We now discuss in more detail the behavior of the different vision models.
First, we trained ConvNets on an MNIST version of symbolic addition for the different training sets. For all tasks, we found that compositional generalization was worse than in-distribution generalization (Fig. 9a) and become worse with smaller distance between the digits (i.e. higher conjunctivity). Further all networks were highly additive ($R^{2}>0.98$). Finally, the values inferred by the network were well consistent with the kernel theory: First, they were an affine function of the true item value for all training sets. Second, extrapolation was worse than interpolation and the task involving extrapolation and interpolation. Notably, interpolation and the task involving both yielded equally good values indicating that these networks may also not behave worse just based on extrapolation vs. interpolation. Further, smaller distance between digits (i.e. more conjunctive inputs) generally exacerbated the value misestimation. Finally, we also considered an asymmetric extrapolation task (“Asymmetric”), which trained on all trials containing an item $[-4]$. Consistent with our theory, we found that this resulted in an affine distortion in values.
Figure 9: Behavior of ConvNets trained on MNIST versions of symbolic addition. a, Mean squared error for different training sets and different evaluation sets. b, Additivity on the different tasks. c, Inferred values.
We then trained a ResNet and a ViT on a CIFAR-10 version of the task. We found that the error on the compositional test dataset was consistently higher than on in-distribution generalization (Fig. 10a). Further, similarly to the ConvNets trained on MNIST, it was particularly high on the asymmetric training set. Notably, the additivity of all of these models was reasonably high ($R^{2}>0.93$, Fig. 10b). Finally, we found that the models all inferred values that were distorted in a roughly affine manner (Fig. 10c).
Figure 10: Behavior of ResNets and ViTs on CIFAR versions of symbolic addition. a, Mean squared error over training. b, Additivity. c, Inferred values for the different tasks.
@@ D.2.4 Coefficient groups
D.2.5 Rich networks
We find that rich networks generalize consistently on CD-1 and CD-2 but not CD-3. They are also perfectly conjunction-wise additive and fail due to a context shortcut (Fig. 11).
@@ D.2.5 Rich networks
D.2.6 Deep neural networks trained on vision data
We found that convolutional neural networks trained on MNIST successfully generalized on CD-1 and CD-2, but not CD-3 (Fig. 12). For smaller distances between digits, the models tended to generalize worse on CD-1 and CD-2 and gradually reverted to chance accuracy (i.e. 0.5) for CD-3. Further, the networks were generally highly additive ($R^{2}>0.95$), but became worse for lower distance (Fig. 12b). Finally, across all distances, they had a high magnitude associated with the context cue, though this magnitude decreased for small distances — consistent with the accuracy of the network increasing from below chance to chance level (Fig. 12c).
@@ D.2.6 Deep neural networks trained on vision data
Finally, we considered ResNets and ViTs trained on a CIFAR version of the task. We found that they generally performed above chance for CD-2 and CD-1, but below chance for CD-3 (Fig. 13a). Further, their additivity was generally high for ResNets ($R^{2}>0.95$) and slightly lower for ViTs ($R^{2}>0.9$) (Fig. 13b) and both networks had a large magnitude associated with the context cue for CD-3, but not CD-2 or CD-1 (Fig. 13c).
Figure 13: Performance of ResNets and ViTs trained on a CIFAR version of context dependence. We plot a, their accuracy on different splits over training, b, the additivity of the networks, and c, the magnitude of the different inferred coefficients.
D.3 Transitive equivalence
We additionally trained ReLU networks of various depth on the transitive equivalence (using $\sigma=0.1$ as initialization magnitude). We consistently found that they were able to generalize to the test set (Fig. 14).
Figure 14: Performance of networks with 100 hidden units per layer and varying depth on transitive equivalence. We used an initialization constant $\sigma=0.1$.
D.4 Invariance and partial exposure
We consider invariance and partial exposure as simple case studies for the memorization leak and shortcut distortion. In both cases, the input consists of two components and the mapping only depends on the first (Fig. 15a). In the invariance case, we don’t see the second component vary at all, in the partial exposure case, we see one instance of the second component. Note that the partial exposure task has previously been studied in the context of network generalization [dasgupta_distinguishing_2022, chan_transformers_2022].
To understand the impact of different representational saliences, we consider the generalization margin $m:=y\hat{y}$ on the test set, where $y$ is the ground-truth label and $\hat{y}$ is the model’s estimate. Because we consider support vector machines, the margin on the training set is one; a smaller margin on the test set indicates worse performance. We determined a mathematical formula for the margins as a function of $\text{Sal}_{2}(1)$. Below we first describe its implications and then how we derived this.
@@ D.4 Invariance and partial exposure
On invariance, we find that the model’s test margin is expressed as $m=\tfrac{\text{Sal}_{2}(1)}{1-\text{Sal}_{2}(1)}$. This means that for a fully compositional representation ($\hat{\kappa}_{1}=0.5$), its training and test margins are both one. However, as $\hat{\kappa}_{1}$ decreases, the model increasingly memorizes the training set, resulting in a decreased margin (Fig. 15b).
@@
On the partial exposure task, if the model used item 1 to solve the task, it would get two out of three training examples correct and could memorize the last data point. This is a statistical shortcut and we find that norm minimization (just as for context dependence) ends up partially relying on this strategy as this decreases the $\ell_{2}$-norm of the readout weights. As a result, the test margin for the partial exposure task decreases even more strongly as a function of $\text{Sal}_{2}(1)$: $m=\tfrac{2\text{Sal}_{2}(1)^{2}}{1-2\text{Sal}_{2}(1)^{2}}$ (where we assume that the similarity between identical trials is one and the similarity between distinct trials is zero) (Fig. 15b).
@@
$m=\frac{(\kappa_{2}+\kappa_{1})(\kappa_{1}-\kappa_{0})}{\kappa_{2}^{2}-\kappa_{1}^{2}}=\frac{\kappa_{1}-\kappa_{0}}{\kappa_{2}-\kappa_{1}}=\frac{(\kappa_{2}-\kappa_{0})\text{Sal}_{2}(1)}{(\kappa_{2}-\kappa_{0})-(\kappa_{1}-\kappa_{0})}=\frac{\text{Sal}_{2}(1)}{1-\text{Sal}_{2}(1)}.$ (50)
@@
$m=\frac{2\text{Sal}_{2}(1)^{2}}{1-2\text{Sal}_{2}(1)^{2}}.$ (54)
D.5 Other mathematical operations
We could consider mathematical operations other than addition as well, considering unobserved assigned values $v_{1}[I_{1}]$ and $v_{2}[I_{2}]$ together with some composition function $C(v_{1}[I_{1}],v_{2}[I_{2}])$. This task will only be additive if the composition function is additive (e.g. if it is subtraction). If it is, e.g. multiplication, division, or exponentiation, the task will be non-additive.
D.6 Logical operations
@@
D.7 Transitive ordering
Transitive ordering is a popular task in cognitive science (often called transitive inference, [mcgonigle_are_1977]). Here the subject is presented with two items $I_{1}$, $I_{2}$ drawn from an unobserved hierarchy $>$. It should then categorize whether $I_{1}>I_{2}$ or $I_{2}>I_{1}$. Crucially, this task can be solved by assigning a rank $r(I_{c})$ to each item and computing the response as $f(I)=r(I_{1})-r(I_{2})$ [lippl_mathematical_2024]. It is therefore additive, in contrast to transitive equivalence. This is also the case if we assume that there are multiple such hierarchies (e.g. $a_{1}>\dotsc,a_{5}$ and $b_{1}>\dotsc,b_{5}$). In this case, the model would generalize to comparisons between these different hierarchies as well.

### v2

@@ B.4 Extended representational analysis
Note that the proof is a minor extension of Lemma S1.3 in Lippl et al. (2024). We present it here in a self-contained manner. We consider a nonlinearity
@@
By prior work (Cho & Saul, 2009; Tsuchida et al., 2018, 2019; Han et al., 2022),
@@ Appendix C Detailed methods
The code required to reproduce all experiments can be found under https://github.com/sflippl/compositional-generalization.
@@ C.1 Models
We fit the kernel models by hand-specifying the kernel and fitting either a support vector regression or classification using scikit-learn Pedregosa et al. (2011). Note that this is equivalent to using a feature basis with the same resulting similarities. However, hand-specifying these similarities enabled us to easily explore the full range of possible representations.
@@
All networks were trained with Pytorch and Pytorch Lightning Paszke et al. (2019). We consider ReLU networks with one hidden layer and $H=1000$ units. We initialize by $\sigma\sqrt{2/H}$, considering $\sigma\in[10^{-6},1]$. In particular, when reporting results on rich networks (without further specification), we assume $\sigma=10^{-6}$. When reporting results on lazy network, we assume $\sigma=1$.
@@
We considered networks with four convolutional layers (kernel size is five, two layers have 32 filters, two have 64 filters) and two densely connected layers (with 512 and 1024 units). Each layer is followed by a ReLU nonlinearity, and the convolutional stage is followed by a max pooling operation. All weights are initialized with He initialization (He et al., 2015).
@@
C.2 Additivity analysis
To analyze how well a conjunction-wise additive computation can describe network behavior, we considered as the set of possible features a concatenation of one-hot vectors coding for each possible conjunction. We then removed all features that are constant at zero on the training dataset and used linear regression to try and predict network behavior on both training and test set for all remaining features. The resulting $R^{2}$ defines the “additivity” of the network behavior (i.e. $R^{2}=1$ indicates full conjunction-wise additivity). Furthermore, we can use the inferred values assigned to these different conjunctions to compare kernel models, feature-learning networks, and vision networks. Note that for the vision networks, we first average the model predictions across all different images instantiating a given compositional input.
Notably, our theory predicts that the constrained conjunction-wise additive function can predict behavior on the test set, but not necessarily on the training set. To test this prediction in vision network, we compared the additivity of the test set alone to the additivity of the entire dataset (Fig. 6). First, we found that the predictivity on the test set was indeed generally very high. Second, we found that it was substantially higher than on the training and test set taken together.
Figure 6: $R^{2}$ of the conjunction-wise additive predictor on the test set as well as on both the training and test set.
Appendix D Compositional task space
@@ D.1 Symbolic addition
Note that the sum over all dual coefficients is given by $2b+c$. Let $p:=|\mathcal{W}|$ and $q:=|\mathcal{V}|-|\mathcal{W}|$. We denote by $\kappa_{c}$ the similarity between inputs overlapping in $c$ components. Then, setting $\delta_{2}:=\kappa_{2}-\kappa_{0}$ and $\delta_{1}:=\kappa_{1}-\kappa_{0}$, the set of dual equations is given by
@@
D.1.2 Behavior of deep ReLU networks
Next, we examined the behavior of deep ReLU networks with varying depth (Fig. 7). On the extrapolation task ($\mathcal{W}=\{0\}$), these networks also had compressed model predictions on the compositional split that were approximately linearly distorted (though for deeper networks, this relationship seemed to have a slight S-shape). Intriguingly, these networks perfectly generalized on the interpolation task ($\mathcal{W}=\{-4,4\}$). For $\mathcal{W}=\{-2,2\}$, they perfectly generalized on the interpolation portion, but not the extrapolation portion.
Figure 7: Model prediction plotted against ground truth for different training sets and depths.
D.1.3 Behavior of vision models
We trained the Convnets on MNIST using seven different training sets $\mathcal{W}$: $\{[0]\}$, $\{[-4],[4]\}$, $\{[-2],[2]\}$, $\{[-4],[0],[4]\}$, $\{[-1],[0],[1]\}$, $\{[-4],[-3],[3],[4]\}$, $\{[-2],[-1],[1],[2]\}$. This allowed us to vary both the size of the training set and whether it involved only interpolation or also extrapolation. We trained these networks for 100 epochs on 20,000 samples. We trained the ResNets and ViTs on CIFAR-10 using the first three training sets listed above. We trained the ResNets for 100 epochs and the ViTs for 200 epochs, training all networks on 40,000 samples. Our findings on these experiments are summarized in Section 6. Further, Fig. 8 depicts, for each training set, the average prediction for each combination of components. We can see that the ConvNet predictions are highly linearly correlated with the ground truth. The ResNet and ViT predictions are also linearly correlated, but exhibit a slightly noisier relationship.
Figure 8: Average prediction for each combination of components plotted against the ground truth for a, ConvNets trained on MNIST, b, ResNets trained on CIFAR-10, and c, ViTs trained on CIFAR-10.
Additionally, we also plot the mean squared error of the networks on the different training sets (Fig. 9). This provides a more conventional (but harder to interpret) measure of performance compared to the slope depicted in Fig. 4. We again see that the mean squared error on the compositional generalization set is decreasing with increasing distance (Fig. 9a). Further, it is decreasing with increasing training set size (Fig. 9), though we note that the differences in the scales of the generalization set for different sets $\mathcal{W}$ renders it harder to directly to compare these values. In contrast, the slopes described in the main text are more easily comparable.
Figure 9: a, Mean squared error as a function of distance. b, Mean squared error as a function of training set size $p$ (different dots corresponding to interpolation versus extrapolation.
@@ D.2.4 Coefficient groups
D.2.5 Feature-learning ReLU networks
We find that feature-learning ReLU networks generalize consistently on CD-1 anREGULARIZATIONd CD-2 but not CD-3. They are also perfectly conjunction-wise additive and fail due to a context shortcut (Fig. 10).
@@ D.2.5 Feature-learning ReLU networks
D.2.6 Vision networks
We additionally analyzed ConvNets for different distances between the digits. We found that convolutional neural networks trained on MNIST successfully generalized on CD-1 and CD-2, but not CD-3 (Fig. 11). For smaller distances between digits, the models tended to generalize worse on CD-1 and CD-2 and gradually reverted to chance accuracy (i.e. 0.5) for CD-3. Further, the networks were generally highly additive ($R^{2}>0.95$), but became worse for lower distance (Fig. 11b). Finally, across all distances, they had a high magnitude associated with the context cue, though this magnitude decreased for small distances — consistent with the accuracy of the network increasing from below chance to chance level (Fig. 11c).
@@ D.2.6 Vision networks
D.3 Invariance and partial exposure
We consider invariance and partial exposure as simple case studies for the memorization leak and shortcut bias. In both cases, the input consists of two components and the mapping only depends on the first (Fig. 12a). In the invariance case, we don’t see the second component vary at all, in the partial exposure case, we see one instance of the second component. Note that the partial exposure task has previously been studied in the context of network generalization (Dasgupta et al., 2022; Chan et al., 2022).
To understand the impact of different representational saliences, we consider the generalization margin $m:=y\hat{y}$ on the test set, where $y$ is the ground-truth label and $\hat{y}$ is the model’s estimate. Because we consider support vector machines, the margin on the training set is one; a smaller margin on the test set indicates worse performance. We determined a mathematical formula for the margins as a function of $S(1;2)$. Below we first describe its implications and then how we derived this.
@@ D.3 Invariance and partial exposure
On invariance, we find that the model’s test margin is expressed as $m=\tfrac{S(1;2)}{1-S(1;2)}$. This means that for a fully compositional representation ($S(1;2)=0.5$), its training and test margins are both one. However, as $S(1;2)$ decreases, the model increasingly memorizes the training set, resulting in a decreased margin (Fig. 12b).
@@
On the partial exposure task, if the model used item 1 to solve the task, it would get two out of three training examples correct and could memorize the last data point. This is a statistical shortcut and we find that norm minimization (just as for context dependence) ends up partially relying on this strategy as this decreases the $\ell_{2}$-norm of the readout weights. As a result, the test margin for the partial exposure task decreases even more strongly as a function of $S(1;2)$: $m=\tfrac{2S(1;2)^{2}}{1-2S(1;2)^{2}}$ (where we assume that the similarity between identical trials is one and the similarity between distinct trials is zero) (Fig. 12b).
@@
$m=\frac{(\kappa_{2}+\kappa_{1})(\kappa_{1}-\kappa_{0})}{\kappa_{2}^{2}-\kappa_{1}^{2}}=\frac{\kappa_{1}-\kappa_{0}}{\kappa_{2}-\kappa_{1}}=\frac{(\kappa_{2}-\kappa_{0})S(1;2)}{(\kappa_{2}-\kappa_{0})-(\kappa_{1}-\kappa_{0})}=\frac{S(1;2)}{1-S(1;2)}.$ (51)
@@
$m=\frac{2S(1;2)^{2}}{1-2S(1;2)^{2}}.$ (55)
D.4 Other mathematical operations
We could consider mathematical operations other than addition as well, considering unobserved assigned values $v_{1}(z_{1})$ and $v_{2}(z_{2})$ together with some composition function $C(v_{1}(z_{1}),v_{2}(z_{2}))$. This task will only be additive if the composition function is additive (e.g. if it is subtraction). If it is, e.g. multiplication, division, or exponentiation, the task will be non-additive.
D.5 Logical operations
@@
D.6 Transitive ordering
Transitive ordering is a popular task in cognitive science (often called transitive inference, (McGonigle & Chalmers, 1977)). Here the subject is presented with two items $z_{1}$, $z_{2}$ drawn from an unobserved hierarchy $>$. It should then categorize whether $z_{1}>z_{2}$ or $z_{2}>z_{1}$. Crucially, this task can be solved by assigning a rank $r(z_{c})$ to each item and computing the response as $f(z)=r(z_{1})-r(z_{2})$ (Lippl et al., 2024). It is therefore additive, in contrast to transitive equivalence. This is also the case if we assume that there are multiple such hierarchies (e.g. $a_{1}>\dotsc,a_{5}$ and $b_{1}>\dotsc,b_{5}$). In this case, the model would generalize to comparisons between these different hierarchies as well.
Appendix E A feature-learning mechanism for non-additive compositional generalization
We proved above that compositionally structured kernel models do not generalize on non-additive tasks, including transitive equivalence. To see if feature learning can overcome this limitation, we trained ReLU networks through backpropagation on transitive equivalence, using a disentangled and uncorrelated input. By varying the initial weight magnitude, we either trained these networks in the kernel/lazy regime or the feature-learning/rich regime. Notably, when trained on symbolic addition and context dependence, the rich networks were well-described by a conjunction-wise additive model (Figs. 7 and 10). On transitive equivalence, however, while the kernel-regime models failed to generalize (as predicted by our theory), the rich networks generalized correctly (Fig. 13a).
To explain why this is the case, we leveraged the insight that rich neural networks are biased to learning weights with a low overall $\ell_{2}$-norm (Lyu & Li (2020); Chizat & Bach (2020); cf. Vardi & Shamir (2021)). In particular, a one-hidden-layer ReLU network tends to learn a sparse set of features Savarese et al. (2019); Chizat & Bach (2020). Transitive equivalence consists of multiple overlapping equality relations (e.g. $A=B\neq D=E$). Notably, ReLU networks such an XOR-type problem by specializing one unit to each conjunction (Brutzkus & Globerson (2019); Saxe et al. (2022); Xu et al. (2023); Fig. 13b). Further, their sparse inductive bias incentivizes ReLU networks to use identical units for overlapping conjunctions (e.g. $(A,B),(A,C)$, and $(B,C)$). This causes the unit to generalize to unseen item combinations (e.g. $(A,C)$), enabling the network to generalize. Importantly, our theoretical argument is corroborated by empirical simulations: each network unit has identical weights for equivalent items (Fig. 13c).
To see whether large-scale neural networks can also benefit from this feature-learning mechanism, we trained ConvNets on an MNIST version of transitive equivalence. The networks were trained for 150 epochs on 20,000 samples. We found that if the digits were presented with a distance of zero, the network did not generalize compositionally at all. However, with increasing distance, the network started to improve its compositional generalization (Fig. 13d), demonstrating that a convolutional network can benefit from this rich compositional motif.